# MoD2T:Model-Data-Driven Motion-Static Object Tracking Method

### v1

@@ V-D Tracking Results
MoD2T’s tracking performance is meticulously evaluated across KITTI, MOT17, and UAVDT datasets, as detailed in Tables IV, V, and VI. The assessment includes key metrics such as HOTA, IDF1, MOTA, and AssA. In the KITTI dataset, MoD2T, synergized with both StrongSORT and DeepOCSORT, exhibits tracking metrics on par with individual methods. Crucially, HOTA, IDF1, MOTA, and AssA scores remain consistent, affirming that MoD2T seamlessly integrates without compromising the intrinsic tracking performance of the base methods. This emphasizes MoD2T’s compatibility, preserving the effectiveness of existing models. The MOT17 dataset reveals a parallel trend, where MoD2T sustains the performance levels of StrongSORT and DeepOCSORT. HOTA, IDF1, MOTA, and AssA scores for MoD2T align closely with standalone methods, corroborating that MoD2T augmentation does not impede the underlying models’ tracking proficiency. UAVDT dataset evaluation further emphasizes MoD2T’s consistent trend in preserving and enhancing tracking performance. Particularly, HOTA, IDF1, MOTA, and AssA scores for MoD2T surpass standalone methods in the AssA metric, suggesting MoD2T’s potential to improve tracking accuracy in specific scenarios. In essence, the results highlight that MoD2T, in collaboration with StrongSORT and DeepOCSORT, maintains the robustness of individual tracking models while substantially enhancing tracking performance. The introduced framework ensures that MoD2T incorporation does not degrade the intrinsic tracking capabilities of the base models. This reinforces MoD2T’s compatibility and efficacy in real-world tracking applications, positioning it as a valuable enhancement for multi-object tracking tasks.
Figure 6: Visualize the sample results of MoD2T on the KITTI test set and some scenarios of MOT17 and UAVDT. The red box represents moving objects, and Teal represents stationary objects. From here, it can be clearly seen that the function of MoD2T to judge the motion status of objects performs very well when the camera is not moving vigorously.
@@ V-E Traditional Method Results
KITTI[50](Camera Fast Moving) 0.013 0.010 0.012 0.009
@@
KITTI[50](Camera Fast Moving) 0.015 0.011 0.015 0.013
@@
Table VII presents the performance of traditional methods on different datasets, with an IoU threshold of 0.5, specifically focusing on correctly tracking moving objects. The results reveal that, at an IoU threshold of 0.5, the performance of traditional methods is generally unsatisfactory across all datasets. Particularly challenging is the KITTI , which involves intense camera motion. When the IoU threshold is relaxed to 0.25, as shown in Table VIII, the proposed method achieves marginally better results in most scenarios. This improvement suggests that, for traditional methods, they can often confirm the presence of a moving object in many cases, but their accuracy in precisely determining the object’s boundaries is limited. However, it becomes evident that when dealing with fast camera motion in the KITTI , the performance of traditional algorithms still falls short. This is also one of the reasons why MoD2T underperforms in such scenarios. Therefore, further research and refinement are necessary to enhance the method’s robustness in fast-paced environments, or alternatively, consider adopting more robust approaches. In summary, traditional methods exhibit limited accuracy when evaluating moving objects at an IoU threshold of 0.5. The proposed method shows some improvement when the IoU threshold is relaxed, but it still faces challenges, particularly in scenarios with fast camera motion, where the performance of traditional algorithms also struggles. To address this issue, continuous efforts should be made to enhance the method’s robustness in fast-motion environments, or explore alternative, more resilient methodologies.
@@ V-F Analysis of Some Parameters
This paper has meticulously expounded on the pivotal fusion parameter configuration, elucidating its intricacies in a preceding section. However, certain parameter settings, such as the parameter denoted as $S_{min}$ in section IV-A, are derived from experiential considerations. To substantiate the rationale behind these parameter choices, supplementary experimental analyses have been conducted. The objective is to furnish a rational explanation for the specified parameter values, shedding light on the nuanced decisions made in their selection within the context of this research endeavor. Regarding $S_{min}$, its configuration in Section IV-D mirrors that of $R$ since both parameters signify the ratio of motion pixels to the entire box pixels. Notably, the ratio $S_{min}/R$ precisely corresponds to the size of bbox. So these two parameters are analyzed together. To illustrate the rationality of our values. We tested the effectiveness of different R values on the sequence of analyzing the rationality of MVF1 design, and the results are shown in the table below.
@@
The choice of R = 0.2 is justified by a careful sensitivity analysis, as depicted in Table IX. A smaller value of R, such as 0.1 or 0.15, makes the MoD2T overly sensitive, leading to the risk of misclassifying static objects as moving. Conversely, a larger R, for instance, 0.3, 0.5, or 0.7, tends to result in reduced sensitivity, potentially causing dynamic objects to be erroneously identified as stationary. Therefore, the selection of R = 0.2 strikes a balance, ensuring optimal sensitivity for motion detection while minimizing the likelihood of false positives or negatives in distinguishing object motion states. In Fig 4, a notable branch exhibits a MOTA score below 0.25. This specific threshold is selected based on the empirical observation that when the MOTA of the traditional method falls below 0.25, it strongly indicates suboptimal performance on sequences of this nature. In such instances, the traditional method proves ineffective. It is essential to clarify that the value of 0.25 serves as a pragmatic threshold, lacking inherent significance beyond being a chosen reference point for assessing traditional method performance on specific sequences.
@@ V-G Limitations
MoD2T primarily excels when dealing with minor camera motion, particularly in scenarios involving rotary-wing unmanned aerial vehicles (UAVs). It effectively distinguishes between stationary and slowly moving objects, resulting in higher MVF1 scores and improved detection capabilities. However, its performance is constrained when the camera experiences significant motion, especially in accurately assessing the motion of mobile objects. This limitation, sensitivity to camera motion, significantly impacts its effectiveness in settings with frequent camera motion, like dynamic outdoor environments or rapid camera movements. Additionally, MoD2T’s performance varies across different object categories, particularly in assessing motion for vehicles and pedestrians, showing significant disparities. This suggests potential biases in specific scenarios. Furthermore, MoD2T’s accuracy in motion assessment is contingent on the quality of the tracking methodology employed. Generally, better tracking methods lead to higher MVF1. MoD2T shines in scenarios with a stationary camera or minor camera motion, proficiently distinguishing stationary from slowly moving objects. However, in contexts with more intense camera motion, especially in assessing mobile object motion, its performance is limited. There are also performance differences across object categories, with better motion assessment for vehicles compared to pedestrians. Future research should focus on addressing these limitations by enhancing MoD2T’s ability to handle pronounced camera motion and improving motion assessment accuracy across all object categories.
@@ VI Conclusion
This paper introduces MoD2T, a novel approach to MOT in video analysis. MoD2T seamlessly integrates traditional mathematical modeling with deep learning-based MOT frameworks, leveraging the strengths of both paradigms. A key innovation is the incorporation of a novel motion-static object tracking algorithm, rooted in motion distance and background features, enhancing accurate determination of object motion status.
MoD2T excels in improving tracking accuracy by dynamically fusing results from traditional methods and deep learning-based MOT frameworks. It intelligently weights and combines tracking outcomes based on their effectiveness, demonstrating superior performance across diverse datasets and scenarios compared to traditional and deep learning methods.
In addition to presenting a novel tracking methodology, we propose MVF1 metric to assess a model’s ability to discern object motion states. We provide a concise analysis of MVF1, emphasizing its significance in comprehensively evaluating tracking performance.
This research contributes fresh perspectives to MOT by showcasing the potential benefits of integrating traditional and deep learning methods. MoD2T not only addresses the limitations of individual methods but also sets the stage for advancements in MOT within video analysis. Future work aims to enhance MoD2T’s performance during high-speed camera movements and explore the feasibility of treating object motion attributes as learnable parameters. This involves leveraging neural networks for prediction and achieving end-to-end motion-based static object discrimination, pushing the boundaries of MOT capabilities in video analysis.
@@
The process of traditional methods in this paper is very similar to [2],which is a mixture of Gaussians, second version method(MOG2). Firstly, we use KIT[57] to correct the camera’s motion. Then, we adopt a dual-mode single Gaussian model with three parameters: age, mean, and variance. The input image is divided into an $N*N$ grid, and each grid employs a single Gaussian model (SGM). This approach effectively reduces computational complexity while maintaining good detection performance. Mathematically, let $G_{i}^{(t)}$ represent the pixel group of grid $i$ at time $t$, $\left|G_{i}^{(t)}\right|$ denotes the number of pixels in this group, and It represents the grayscale value of pixel $j$ at time $t$. The update functions for mean, variance, and age in the single Gaussian model are given by (18):
@@
We represent the parameters of the candidate model and the current model for grid $i$ at time $t$ as $\mu_{B,i}^{(t)},\sigma_{B,i}^{(t)}$and $\alpha_{B,i}^{(t)}$, respectively.If the observed difference between the average value $M_{i}^{(t)}$and the squared value $\mu_{B,i}^{(t)}$, is less than the critical value related to $\alpha_{B,i}^{(t)}$, as shown in (22):
@@
If neither of the two conditions mentioned above is met, the candidate background model is updated with the current observed value. At the same time, if (23) is satisfied:
@@
i.e., when the age of the candidate background model is greater than the current background model’s age, the two models will exchange their roles. Finally, after determining the models, for the j-th pixel within the i-th pixel set, if (25) is satisfied, the pixel is recognized as a foreground pixel, where d is a threshold value. Note that only the current background model is used to decide the foreground.

### v2

@@ V-C Motion State Judgment Results
Directing attention to the KITTI dataset, encompassing scenes characterized by diverse camera motions, MoD2T demonstrates noteworthy enhancements in the MVF1 score, particularly in scenarios involving slow camera motion or a stationary camera. The method excels in accurately discerning static from moving objects. However, despite these advancements, the overall tracking performance across all methods remains consistent, indicating that the improvements predominantly contribute to the discrimination of static or slowly moving objects, rather than constituting a substantial enhancement to tracking capabilities. Notably, MoD2T’s capacity to discern object motion diminishes in scenarios characterized by rapid camera motion.
In the UAVDT dataset, MoD2T exhibits a significant elevation in the MVF1 score compared to deep learning-based multiple object tracking (MOT) methods, such as StrongSORT and DeepOCSORT. Analogous to the KITTI dataset, the tracking performance remains relatively unaffected, underscoring MoD2T’s proficiency in detecting static or slowly moving objects without substantial gains in overall tracking capabilities.
Similarly, in the MOT17 dataset, analogous trends emerge. MoD2T’s efficacy in determining object motion is notably superior in scenarios with slow camera motion or a stationary camera (e.g., MOT17-03), in contrast to situations featuring rapid camera motion (e.g., MOT17-07). Moreover, when cross-referencing datasets, such as UAVDT and KITTI, wherein camera motion is minimal, it becomes apparent that discerning human motion is considerably more challenging than detecting vehicle movement.
[…]
This alignment in outcomes is likely attributable to shared foundational principles or algorithmic components, resulting in similar responses to diverse motion state scenarios. Both methods may employ comparable strategies to address challenges posed by stationary or slowly moving objects versus swiftly moving entities.
Minor disparities observed in specific instances may stem from dataset-specific attributes or variations in parameters, such as the tuning parameter $\lambda$ in motion state judgment module. Nevertheless, the overall trend of closely aligned performance suggests a robust and consistent behavior exhibited by both tracking approaches across diverse scenarios.
In summary, the consistent alignment of results across various datasets emphasizes the comparable effectiveness of DeepOCSORT and StrongSORT in motion state judgment. Despite potential differences in their underlying architectures, both methods consistently deliver similar tracking outcomes, showcasing their proficiency in addressing diverse motion scenarios encountered in real-world tracking applications.
@@ V-D Tracking Results
MoD2T’s tracking performance is meticulously evaluated across KITTI, MOT17, and UAVDT datasets, as detailed in Tables IV, V, and VI. The assessment includes key metrics such as HOTA, IDF1, MOTA, and AssA.
In the KITTI, MoD2T, synergized with both StrongSORT and DeepOCSORT, exhibits tracking metrics on par with individual methods. Crucially, HOTA, IDF1, MOTA, and AssA scores remain consistent, affirming that MoD2T seamlessly integrates without compromising the intrinsic tracking performance of the base methods. This emphasizes MoD2T’s compatibility, preserving the effectiveness of existing models.
The MOT17 reveals a parallel trend, where MoD2T sustains the performance levels of StrongSORT and DeepOCSORT. HOTA, IDF1, MOTA, and AssA scores for MoD2T align closely with standalone methods, corroborating that MoD2T augmentation does not impede the underlying models’ tracking proficiency.
UAVDT dataset evaluation further emphasizes MoD2T’s consistent trend in preserving and enhancing tracking performance. Particularly, HOTA, IDF1, MOTA, and AssA scores for MoD2T surpass standalone methods in the AssA metric, suggesting MoD2T’s potential to improve tracking accuracy in specific scenarios.
In essence, the results highlight that MoD2T, in collaboration with StrongSORT and DeepOCSORT, maintains the robustness of individual tracking models while substantially enhancing tracking performance. The introduced framework ensures that MoD2T incorporation does not degrade the intrinsic tracking capabilities of the base models. This reinforces MoD2T’s compatibility and efficacy in real-world tracking applications, positioning it as a valuable enhancement for multi-object tracking tasks.
Figure 6: Visualize the sample results of MoD2T on the KITTI and some scenarios of MOT17 and UAVDT. The red box represents moving objects, and teal represents stationary objects. From here, it can be clearly seen that the function of MoD2T to judge the motion status of objects performs very well when the camera is not moving vigorously.
@@ V-E Traditional Method Results
KITTI[55](Moving Fast) 0.013 0.010 0.012 0.009
@@
KITTI[55](Moving Fast) 0.015 0.011 0.015 0.013
@@
Table VII presents the performance of traditional methods on different datasets, with an IoU threshold of 0.5, specifically focusing on correctly tracking moving objects. The results reveal that, at an IoU threshold of 0.5, the performance of traditional methods is generally unsatisfactory across all datasets. Particularly challenging is the KITTI , which involves intense camera motion.
When the IoU threshold is relaxed to 0.25, as shown in Table VIII, the proposed method achieves marginally better results in most scenarios. This improvement suggests that, for traditional methods, they can often confirm the presence of a moving object in many cases, but their accuracy in precisely determining the object’s boundaries is limited.
However, it becomes evident that when dealing with fast camera motion in the KITTI , the performance of traditional algorithms still falls short. This is also one of the reasons why MoD2T underperforms in such scenarios. Therefore, further research and refinement are necessary to enhance the method’s robustness in fast-paced environments, or alternatively, consider adopting more robust approaches.
In summary, traditional methods exhibit limited accuracy when evaluating moving objects at an IoU threshold of 0.5. The proposed method shows some improvement when the IoU threshold is relaxed, but it still faces challenges, particularly in scenarios with fast camera motion, where the performance of traditional algorithms also struggles. To address this issue, continuous efforts should be made to enhance the method’s robustness in fast-motion environments, or explore alternative, more resilient methodologies. Meanwhile, In comparison to Tables IV, V, and VI, MoD2T demonstrates significant advantages in tracking performance over traditional method.
@@ V-F Analysis of Some Parameters
This paper has meticulously expounded on the pivotal fusion parameter configuration, elucidating its intricacies in a preceding section. However, certain parameter settings, such as the parameter denoted as $S_{min}$ in section IV-A, are derived from experiential considerations. To substantiate the rationale behind these parameter choices, supplementary experimental analyses have been conducted. The objective is to furnish a rational explanation for the specified parameter values, shedding light on the nuanced decisions made in their selection within the context of this research endeavor.
Regarding $S_{min}$, its configuration in Section IV-C mirrors that of $R$ since both parameters signify the ratio of motion pixels to the entire box pixels. Notably, the ratio $S_{min}/R$ precisely corresponds to the size of bbox. So these two parameters are analyzed together.
To illustrate the rationality of our values. We tested the effectiveness of different R values on the sequence of analyzing the rationality of MVF1 design, and the results are shown in the table below.
@@
The choice of R = 0.2 is justified by a careful sensitivity analysis, as depicted in Table IX. A smaller value of R, such as 0.1 or 0.15, makes the MoD2T overly sensitive, leading to the risk of misclassifying static objects as moving. Conversely, a larger R, for instance, 0.3, 0.5, or 0.7, tends to result in reduced sensitivity, potentially causing dynamic objects to be erroneously identified as stationary. Therefore, the selection of R = 0.2 strikes a balance, ensuring optimal sensitivity for motion detection while minimizing the likelihood of false positives or negatives in distinguishing object motion states.
In Figure 3, a notable branch exhibits a MOTA score below 0.25. This specific threshold is selected based on the empirical observation that when the MOTA of the traditional method falls below 0.25, it strongly indicates suboptimal performance on sequences of this nature. In such instances, the traditional method proves ineffective. It is essential to clarify that the value of 0.25 serves as a pragmatic threshold, lacking inherent significance beyond being a chosen reference point for assessing traditional method performance on specific sequences.
@@ V-G Limitations
MoD2T primarily excels when dealing with minor camera motion, particularly in scenarios involving rotary-wing UAVs. It effectively distinguishes between stationary and slowly moving objects, resulting in higher MVF1 scores and improved detection capabilities. However, its performance is constrained when the camera experiences significant motion, especially in accurately assessing the motion of mobile objects. This limitation, sensitivity to camera motion, significantly impacts its effectiveness in settings with frequent camera motion, like dynamic outdoor environments or rapid camera movements. Additionally, MoD2T’s performance varies across different object categories, particularly in assessing motion for vehicles and pedestrians, showing significant disparities. This suggests potential biases in specific scenarios.
Furthermore, MoD2T’s accuracy in motion assessment is contingent on the quality of the tracking methodology employed. Generally, better tracking methods lead to higher MVF1. MoD2T shines in scenarios with a stationary camera or minor camera motion, proficiently distinguishing stationary from slowly moving objects. However, in contexts with more intense camera motion, especially in assessing mobile object motion, its performance is limited. There are also performance differences across object categories, with better motion assessment for vehicles compared to pedestrians. Future research should focus on addressing these limitations by enhancing MoD2T’s ability to handle pronounced camera motion and improving motion assessment accuracy across all object categories.
@@ VI Conclusion
In this paper, we introduce a novel benchmark in the domain of MOT: motion state. Subsequently, addressing this benchmark, we present MoD2T, a pioneering approach to MOT in video analysis. MoD2T seamlessly integrates traditional mathematical modeling with deep learning-based MOT frameworks, harnessing the advantages of both paradigms. A key innovation lies in the incorporation of a novel motion-static object tracking algorithm, grounded in motion distance and background features, thereby enhancing the precise determination of object motion status.
Furthermore, we propose the MVF1 metric to assess a model’s ability to discern object motion states. We provide a concise analysis of MVF1, emphasizing its significance in comprehensively evaluating tracking performance.
This paper presents a fresh perspective in the field of multiple object tracking (MOT) by demonstrating the potential benefits of integrating traditional and deep learning methods. MoD2T not only addresses the limitations of various methods but also lays the groundwork for advancements in MOT within video analysis, owing to its remarkable capability in discerning object motion states. In future work, we intend to explore the integration of 3D object detection models. We believe that this approach could alleviate performance issues observed under rapid camera movement, particularly from the bird’s-eye view (BEV) perspective.
@@
The process of traditional methods in this paper is very similar to [2],which is a mixture of Gaussians, second version method(MOG2).
Firstly, we use KIT[62] to correct the camera’s motion.
Then, we adopt a dual-mode single Gaussian model with three parameters: age, mean, and variance. The input image is divided into an $N*N$ grid, and each grid employs a single Gaussian model (SGM). This approach effectively reduces computational complexity while maintaining good detection performance.
Mathematically, let $G_{i}^{(t)}$ represent the pixel group of grid $i$ at time $t$,$\left|G_{i}^{(t)}\right|$ denotes the number of pixels in this group, and It represents the grayscale value of pixel $j$ at time $t$. The update functions for mean, variance, and age in the single Gaussian model are given by (17):
@@
We represent the parameters of the candidate model and the current model for grid $i$ at time $t$ as $\mu_{B,i}^{(t)},\sigma_{B,i}^{(t)}$and $\alpha_{B,i}^{(t)}$, respectively.If the observed difference between the average value $M_{i}^{(t)}$and the squared value $\mu_{B,i}^{(t)}$, is less than the critical value related to $\alpha_{B,i}^{(t)}$, as shown in 21:
@@
If neither of the two conditions mentioned above is met, the candidate background model is updated with the current observed value.
At the same time, if (23) is satisfied:
@@
i.e., when the age of the candidate background model is greater than the current background model’s age, the two models will exchange their roles.
Finally, after determining the models, for the j-th pixel within the i-th pixel set, if (24) is satisfied, the pixel is recognized as a foreground pixel, where d is a threshold value. Note that only the current background model is used to decide the foreground.